Which represents the exponential form of log5(1/5)= -1


a. 1/5 5 = -1

b. 1/5-1= 5

c. 51/5 = -1

d. 5-1= 1/5



1
Expert's answer
2021-05-13T13:31:43-0400


Given

"log_5(\frac{1}{5})=-1\\\frac{1}{5}=5^{-1}\\5^{-1}=\frac{1}{5}"


option d is correct